OpenStudy (anonymous):

6y-2x=20 find y plz help

OpenStudy (scorcher219396):

Assuming you mean "isolate y" or "solve for y", Put anything that has a y on one side, and everything else on the other 6y-2x=20 +2x +2x 6y=20+2x Then divide everything by the coefficient of the y (in this case, 6) 6y=20+2x /6 /6 /6 y=10/3 + (1/3)x